DAFNE Trouble Ticket

Home
Show
Open
Update
Close
Forward
Acknowledge
Status
Report
Logout

Week Program & Report

Week N. 21 (19/05/20 / 19/05/27), Run Coordinator on duty: Boscolo Manuela

 

Monday 20 May

MD on both beams. Goal is  to increase the   beams current and optimize the beams trajectories.

Measurement of betatron tunes:  (Qx(e-) = 0.0726)

 

Tuesday 21 May

 Opening scheduled at ~ 10 am.

RAI TV video shooting from ~10 am to  ~15.30.

Aligment measurements performed in the new IR area.

Other maintanence activities in parallel, some listed below:

Water leakage on one wiggler detected and repaired.

Vacuum pump in the new IR area substituted.

Scrapers opening jaws  and vacuum valves  rechecked.

Work on the Siddharta detector.

Diagnostics: all BPM of IR1 checked with reflectometry method. BPMPS100: one of the 4 cables was disconnected, now ok.

 

15:30 About ready to restart the operation.

 

 Wednesday 22 May

Optics optimization on both beams.

e- beam : tunes measurement, optics bumps , sextupoles and octupoles .

                 80 bunches with ~56mA 

 

 

Last update: 2019-05-22 16:06:40

User info: Anonymous on
107.23.129.77
Author: Accelerator TT Group Developpers: D. Maselli & G. Mazzitelli